# Exact output tracking for the one-dimensional heat equation and applications to the interpolation problem in Gevrey classes of order 2

### v1

@@ 1.1 Output tracking
Concerning the exact tracking of (1), only a few results already exist in the literature, each in the case of finite $T$. On the one hand, any $y\in\mathcal{Y}(0,T)$ satisfies $y\in C^{\infty}[0,T]$ together with the two conditions
@@
Such a regularity result can be deduced from Gevrey’s estimates [gevrey]; see also [benedetto, Proposition 12.1, Chapter 5]. Moreover, in the above, the parameter $R$ is uniform with respect to $T$ and $y$. On the other hand, it has been established by Martin Rosier and Rouchon in [reachable_martin, Proposition 3.1] (see also Hölmgren [holmgren] for connected results) using the flat-output method that a sufficient condition for a signal $\varphi\in C^{\infty}[0,T]$ to belong to $\mathcal{Y}(0,T)$ is that there is some $R>1$ such that $\varphi$ satisfies (2) together with (3).
@@
hence, regarding $f$ as a function of two real variables we have $\nabla f\in L^{2}(\Omega)$. Therefore, by standard procedures (see e.g. [mazya, §1.1.11]), we deduce that $f$ belongs to the Sobolev space $W^{1,2}(\Omega)$ [^1] We do not use the notation $H^{1}(\Omega)$ for the Sobolev space $W^{1,2}(\Omega)$ in order to save the letter $H$ for the Hardy space $H^{2}$., where $\Omega$ is seen as a subset of $\mathbb{R}^{2}$. Thus,
$\mathcal{R}=\{f\in\operatorname{Hol}(\Omega)\cap W^{1,2}(\Omega),\,f\mbox{ even}\},$ (6)
where $\operatorname{Hol}(D)$ stands for the set of holomorphic functions in $D$, whenever $D$ is an open subset of $\mathbb{C}$.
Now take a solution $z$ of (1). It is in $C^{\infty}((0,T)\times(0,1))$, and can be extended as an even and $C^{\infty}$ function on $(0,T)\times(-1,1)$, which solves the heat equation therein. The flat-output method is a general procedure to parameterize the state of a system as a power series of the successive time derivatives of a well-chosen output, we refer to [fliess] for a general overview. For the heat equation (1), this method has been implemented by developing $z$ as power series of $x$, we refer to [holmgren, laroche, reachable_martin]. Notably, for any $u\in L^{2}(0,T)$, the solution $z$ of (1) writes
@@
From (7) and Cauchy’s inequalities, we see that the largest $R$ such that $y\in G^{2,R}[0,T]$ is also the largest $R$ such that $z(t,x)$ is holomorphic with respect to $x$ in $D(0,R)$. From (6), it follows that the largest such $R$ is $R=1/\sqrt{2}$, therefore, the trackable functions are Gevrey of order 2 and radius $R=1/\sqrt{2}$, and no better.
@@
Let $0<T<\infty$ and $\varphi\in C^{\infty}_{(0)}[0,T]$ satisfying (8). Then for every $0<\delta<T$, we have $\varphi\in\mathcal{Y}(0,T-\delta)$.
@@ 1.2 Some intermediate results of interest
where $\hat{\penalty 10000\ }$ stands for the Laplace transform and $\mathbb{C}_{+}$ is the open right half-plane (see §2.4). This can be rewritten as
@@
where $\hat{u}$ belongs to the Hardy space $H^{2}(\mathbb{C}_{+})$ and $1/\phi\in H^{\infty}(\mathbb{C}_{+})$, hence $\hat{\dot{y}}\in H^{2}(\mathbb{C}_{+})$ (see §2.1 for the definition of Hardy spaces). From (12) we see that the description of $\mathcal{Y}(0,\infty)$ is equivalent, through the Laplace transform, to the description of these $F\in H^{2}(\mathbb{C}_{+})$ such that $\phi F$ lies again in $H^{2}(\mathbb{C}_{+})$. Proposition 6 below allows us to characterize such functions $F$, under a growth assumption on $\phi$.
@@
Let $\phi$ be an entire function of order <1<1222see Definition 18. Then, for every $F\in H^{2}(\mathbb{C}_{+})$, the product $\phi F$ lies in $H^{2}(\mathbb{C}_{+})$ if and only if
@@
Observe that for $\phi$ an entire function of order <1<1 and $F\in H^{2}(\mathbb{C}_{+})$, the product $\phi F$ automatically has the non-tangential boundary values $\phi(i\xi)F(i\xi)$ (see §2.1). Thus, $\phi F$ lies in $H^{2}(\mathbb{C}_{+})$ if and only if the two following conditions are satisfied
$\phi(i\cdot)F(i\cdot)\in L^{2}(\mathbb{R}),\quad\mbox{and}\quad\operatorname{supp}\mathcal{F}[\phi(i\cdot)F(i\cdot)]\subset(-\infty,0],$ (13)
where $\mathcal{F}$ is the Fourier transform. Proposition 6 thus states that for $\phi$ of order <1<1, the first condition of (13) implies the second one. This can be seen with the following heuristic computation
$\operatorname{supp}\mathcal{F}[\phi(i\cdot)F(i\cdot)]$ $=\operatorname{supp}\mathcal{F}[\phi(i\cdot)]*\mathcal{F}[F(i\cdot)]$
$\subset(\operatorname{supp}\mathcal{F}[\phi(i\cdot)])+(\operatorname{supp}\mathcal{F}[F(i\cdot)])$
$\subset\{0\}+(-\infty,0]\subset(-\infty,0],$
where we have used $\operatorname{supp}\mathcal{F}[\phi(i\cdot)]\subset\{0\}$. Unfortunately, the above computation remains heuristic as almost every the written calculations hardly make sense. For instance, a function $\phi$ entire on $\mathbb{C}$ with order <1<1 may not even lie in $\mathcal{S}^{\prime}(\mathbb{R})$ (e.g. $\phi(s)=\cosh\sqrt{s}$), hence the definition of $\mathcal{F}\phi$ is tedious. However, the outcome of the above heuristic
$\operatorname{supp}\mathcal{F}[\phi(i\cdot)F(i\cdot)]\subset(-\infty,0],$
will be shown to be true (see Lemma 19). This result will allow us to give a meaning to the assertion $\operatorname{supp}\mathcal{F}[\phi(i\cdot)]\subset\{0\}$, and to use it in order to prove Proposition 6. The latter assertion should be compared with the Paley–Wiener theorem [rudin, Theorem 19.3, p. 370], asserting that if $\phi\in\operatorname{Hol}(\mathbb{C})\cap L^{2}(\mathbb{R})$ is of order $1$ and satisfies
@@
for some $A>0$, then $\operatorname{supp}\mathcal{F}\phi\subset[-A,A]$.
@@
$\|\varphi\|_{\mathcal{G}_{s,R,\gamma}}^{2}:=\sum_{n=0}^{\infty}\left(\frac{\|\varphi^{(n)}\|_{L^{2}(\mathbb{R})}}{M_{n}}\right)^{2}<\infty,\quad M_{n}:=\left(\frac{ns}{eR}\right)^{ns}(1+n)^{-s\gamma+1/4}.$ (15)
The above defines a Hilbert space $\mathcal{G}_{s,R,\gamma}$, which is continuously and densely embedded in $L^{2}(\mathbb{R},dx)$.
@@
To our knowledge, Theorem 7 is new even for $\gamma=0$. Indeed, it is well-established is that a function $\varphi\in C_{c}^{\infty}(\mathbb{R})$ such that
$\exists A,C>0,\quad\forall k\in\mathbb{N},\quad\forall x\in\mathbb{R},\quad|\varphi^{(k)}(x)|\leq CA^{k}k^{ks},$
@@
$\exists B,C^{\prime}>0,\quad\forall\xi\in\mathbb{R},\quad|\mathcal{F}\varphi(\xi)|\leq C^{\prime}e^{-B|\xi|^{1/s}},$
and the converse also holds (see e.g. [gs2, Chapter IV],[Rodino, Theorem 1.6.1]). In the previous works we are aware of, a link between $A$ and $B$ can be made explicit, but it is never sharp.
@@ 1.3 Consequences for the interpolation problem in Gevrey-$2$ class
Consider a sequence of positive numbers $(m_{k})$, possibly taking the value $\infty$, and the associated Denjoy-Carleman class
$C^{\infty}(m_{k}):=\left\{\varphi\in C^{\infty}(\mathbb{R}):\exists C,A>0,\quad\forall k\in\mathbb{N},\quad\forall x\in\mathbb{R},\quad|\varphi^{(k)}(x)|<CA^{k}m_{k}\right\}.$
Such classes naturally arise in several contexts, such as the regularity theory of PDEs [holmgren, temam], the asymptotic developments of the solutions of algebraic ODEs [ramis, devissage], or the control theory of parabolic PDEs [reachable_martin, rosier, laurent]. We refer to [bilodeau] for a historical survey. Some interesting properties of the class $C^{\infty}(m_{k})$ can be described by the so-called Borel operator
$\mathcal{B}:C^{\infty}(m_{k})\rightarrow\mathbb{C}^{\mathbb{N}},\quad\mathcal{B}\varphi:=(\varphi^{(k)}(0))_{k=0}^{\infty}.$
A theorem due to Denjoy and Carleman gives a necessary and sufficient condition on the sequence $(m_{k})$ for the operator $\mathcal{B}$ to be injective. If this is the case, the class $C^{\infty}(m_{k})$ is called quasi-analytic, as it shares the following feature with analytic functions: a function $\varphi\in C^{\infty}(m_{k})$ identically vanishes as soon as every its derivatives at 0 vanish. In the opposite case, we call $C^{\infty}(m_{k})$ a non quasi-analytic class.
Another important property of $C^{\infty}(m_{k})$ is the characterization of the range of $\mathcal{B}$, which we will refer to as the interpolation problem. Two extreme cases must be mentioned. The first one is when $\mathcal{B}$ acts on $C^{\infty}(\mathbb{R})$ (i.e. $m_{k}\equiv\infty$). In this case, the operator $\mathcal{B}$ is surjective, which is commonly referred to as Borel’s Theorem [^3] Borel showed this result in [Bobo], although it seems to have been proved first by Peano, see [borel]. The other extreme case is when $\mathcal{B}$ acts on $\operatorname{Hol}(D(0,R))$. In this case the Cauchy-Hadamard formula yields that a sequence $(a_{k})$ belongs to $\mathcal{B}\operatorname{Hol}(D(0,R))$ if and only if
$\forall 0<\epsilon<R,\quad\exists c>0,\quad\forall k\in\mathbb{N},\quad|a_{k}|\leq c\frac{k!}{(R-\epsilon)^{k}}.$
The case of $\mathcal{B}$ acting on a non quasi-analytic class can thus be viewed as intermediate between these two extreme cases. At this point it is meaningful to make the following observation: if the class $C^{\infty}(m_{k})$ is non quasi-analytic, then one can build a cutoff function belonging to the class $C^{\infty}(m_{k})$, and consequently the Borel operator is local:
$\forall\epsilon>0,\quad\mathcal{B}C^{\infty}(m_{k})=\mathcal{B}\{\varphi\in C^{\infty}(m_{k}):\operatorname{supp}\varphi\subset(-\epsilon,+\epsilon)\}.$
Many works have addressed the description of the range of $\mathcal{B}$ over $C^{\infty}(m_{k})$, for general sequences $(m_{k})$, we refer to [carleson, petzsche, kiro, wahde, thilliez, carleman]. From the previously cited papers, we know that under suitable assumptions on $(m_{k})$ we have
$\mathcal{B}C^{\infty}(m_{k})=\{(a_{k})\in\mathbb{C}^{\mathbb{N}}:\exists C,A>0,\quad\forall k\in\mathbb{N},\quad|a_{k}|\leq CA^{k}m_{k}\}.$
We understand this result as: given a sequence $(a_{k})_{k\in\mathbb{N}}$ and $A>0$ such that $|a_{k}|\lesssim A^{k}m_{k}$, we can find $B\geq A$ and $\varphi\in C^{\infty}(\mathbb{R})$ such that $\varphi^{(k)}(0)=a_{k}$ and $|\varphi^{(k)}(x)|\lesssim B^{k}m_{k}$, for every $k$ and $x$. Passing from $A$ to $B$ can be thought of as relaxing the imposed growth, which amounts, in a sense, to losing a bit of regularity. We say that the function $\varphi$ interpolates the sequences $(a_{k})$ (at the point 0).
For the applications in controllability theory we have in mind [reachable_martin, rosier, laurent], understanding how $B$ depends on $A$ is crucial in order to obtain sharp reachability results. From now on we restrict ourselves to non quasi-analytic Gevrey classes, that is we take $m_{k}=k^{ks}$ for some $s>1$  [^4] In this part, in order to be in accordance with the existing literature and notably [mitjagin], we chose to state our results by using the usual definition of Gevrey classes in terms of powers rather than in terms of factorials, as in the previous section.. We consider the set $C^{\infty}(k^{ks},A)$ of functions $\varphi\in C^{\infty}(\mathbb{R})$ such that $|\varphi^{(k)}(x)|\lesssim A^{k}k^{ks}$ for fixed $A>0$. A trivial necessary condition for a sequence $(a_{k})$ to be interpolated by a function in $C^{\infty}(k^{ks},A)$ is that it satisfies the Gevrey type estimate $|a_{k}|\lesssim A^{k}k^{ks}$. However, this turns out to be far from sufficient, as shown in the following fact, due to Mitjagin [mitjagin]: for every $\epsilon,A>0$, a sufficient condition for the interpolation of a sequence $(a_{k})$ is given by
$|a_{k}|\lesssim A^{k}k^{ks}\Longrightarrow(a_{k})\in\mathcal{B}C^{\infty}(k^{ks},\Gamma_{s}A+\epsilon).$ (16)
where the constant $\Gamma_{s}>1$ is defined by $\Gamma_{s}:=\cos^{-s}(\pi/2s)$. The above condition is sharp in the sense that
$\forall A,\epsilon>0,\quad\exists(a_{k})\in\mathbb{C}^{\mathbb{N}},\quad|a_{k}|\lesssim A^{k}k^{ks}\mbox{ and }(a_{k})\notin\mathcal{B}C^{\infty}(k^{ks},\Gamma_{s}A-\epsilon).$ (17)
The paper [mitjagin] is written in Russian, and its translation [mitjagin_english] is rather hard to access. Furthermore, in view of its size (4 pages), the published version seems to offer only a sketch of the proof rather than a fully detailed demonstration. Hence, for the sake of completeness, we give in Appendix A a detailed proof of Theorems 1 and 1a of [mitjagin].
As a consequence of our study of the trackable subspace $\mathcal{Y}(0,T)$, for an arbitrary sequence $(a_{k})$ we give the optimal (smallest) $B>0$ such that $(a_{k})\in\mathcal{B}C^{\infty}(k^{ks},B)$, when $s=2$. For a sequence $(a_{k})$ we define the number
$R_{a}:=\sup\{R>0:\exists\varphi\in\mathcal{G}_{2,R,0},\quad\forall k\in\mathbb{N},\quad\varphi^{(k)}(0)=a_{k}\},$
where it is understood that $R_{a}=0$ if the above supremum runs over the empty set. We point out that in the above definition, there is no loss of generality in assuming that $\varphi$ has compact support, or to replace the space $\mathcal{G}_{2,R,0}$ by $G^{2,R}$. For instance, one can show that
$R_{a}=\sup\{R>0:\exists\epsilon>0,\quad\exists\varphi\in G^{2,R}[-\epsilon,\epsilon],\quad\forall k\in\mathbb{N},\quad(\varphi^{(k)}(-\epsilon),\varphi^{(k)}(0),\varphi^{(k)}(\epsilon))=(0,a_{k},0)\},$
where $G^{2,R}[-\epsilon,\epsilon]$ is defined Similarly to in (3), replacing the interval $[0,T]$ by $\epsilon,\epsilon]$.
@@
The notation $A^{2}(\Omega)$ stands for the Bergman space on $\Omega$, i.e. the space of functions on $\Omega$ that are holomorphic and square-integrable on $\Omega$. By abuse of notation, we shall say that a power series $\sum a_{n}\zeta^{n}$ belongs to $A^{2}(\Omega)$ if it has radius of convergence $\geq 1/\sqrt{2}$, and the sum function has an extension to $\Omega$ which is holomorphic and square-integrable.
Corollary 9 is a direction consequence of the following proposition, which characterizes the range of the Borel operator over some Hilbert spaces of Gevrey functions. We will use the following notation: for an interval $I\subset\mathbb{R}$ and parameters $s,R>0$ and $\gamma\in\mathbb{R}$, we let $\mathcal{G}_{s,R,p}(I)$ be the Hilbert space defined similarly as in (15), with $L^{2}(I)$ in place of $L^{2}(\mathbb{R})$.
Proposition 10.
Let $I\subset\mathbb{R}$ be a non trivial interval containing 0 in its interior, $R>0$ and $p\in\mathbb{Z}$. Then
and
The rest of the paper is organized as follows. In Section 2 we show our tracking results Theorems 1 and 2, as well as the intermediate results Proposition 6 and Theorem 7, and finally Proposition 4. In Section 3 we explain how our methods can be applied to systems close to (1), having different boundary conditions or smoother control laws, or where the output is given by the Dirichlet trace at a given point of $(0,1)$. In Section 4 we show Proposition 10. In Appendix A we study the interpolation problem in Gevrey classes, notably with a detailed proof of Mitjagin’s quoted results (16) and (17).

### v2

@@ 1.1 Output tracking
Concerning the exact tracking of (1), only few results already exist in the literature, which are all in the case of finite $T$. On the one hand, any $y\in\mathcal{Y}(0,T)$ satisfies $y\in C^{\infty}[0,T]$ together with the two conditions
@@
Such a regularity result can be deduced from Gevrey’s estimates [benedetto, Proposition 12.1, Chapter 5]. Moreover, in the above, the parameter $R$ is uniform with respect to $T$ and $y$. On the other hand, it has been established by Martin Rosier and Rouchon in [reachable_martin, Proposition 3.1] (see also Hölmgren [holmgren] for connected results) using the flat-output method that a sufficient condition for a signal $\varphi\in C^{\infty}[0,T]$ to belong to $\mathcal{Y}(0,T)$ is that there is some $R>1$ such that $\varphi$ satisfies (2) together with (3).
@@
hence, regarding $f$ as a function of two real variables we have $\nabla f\in L^{2}(\Omega)$. Therefore, by standard procedures (see e.g. [mazya, §1.1.11]), we deduce that $f$ belongs to the Sobolev space $H^{1}(\Omega)$, where $\Omega$ is seen as a subset of $\mathbb{R}^{2}$. Thus,
$\mathcal{R}=\{f\in\mathcal{H}(\Omega)\cap H^{1}(\Omega),\,f\mbox{ even}\},$ (6)
where $\mathcal{H}(D)$ stands for the set of holomorphic functions on $D$, whenever $D$ is an open subset of $\mathbb{C}$.
Now take a solution $z$ of (1). It lies in $C^{\infty}((0,T)\times(0,1))$, and can be extended as an even function of $x$, lying in $C^{\infty}((0,T)\times(-1,1))$, which solves the heat equation on $(0,T)\times(-1,1)$. The flat-output method is a general procedure to parameterize the state of a system as a power series of the successive time derivatives of a well-chosen output, we refer to [fliess] for a general overview. For the heat equation (1), this method has been implemented by developing $z$ as power series of $x$, we refer to [holmgren, laroche, reachable_martin]. Notably, for any $u\in L^{2}(0,T)$, the solution $z$ of (1) writes
@@
From (7) and Cauchy’s inequalities, we see that the largest $R$ such that $y\in G^{2,R}[0,T]$ is also the largest $R$ such that $z(t,x)$ is holomorphic with respect to $x$ in $D(0,R)$. From (6), it follows that the largest such $R$ is $R=1/\sqrt{2}$, therefore, trackable functions are Gevrey of order 2 and radius $R\leq 1/\sqrt{2}$, and no better in general. Theorem 1 makes this more precise and more importantly provides a converse statement.
@@
Let $0<T<\infty$ and $\varphi\in C^{\infty}_{(0)}[0,T]$ satisfying (8) (where $y$ is replaced by $\varphi$). Then for every $0<\delta<T$, we have $\varphi\in\mathcal{Y}(0,T-\delta)$.
@@ 1.2 Some intermediate results of interest
where $\hat{\penalty 10000\ }$ stands for the Laplace transform and $\mathbb{C}_{+}$ is the open right half-plane. This can be rewritten as
@@
where $\hat{u}$ belongs to the Hardy space $\mathcal{H}^{2}(\mathbb{C}_{+})$ and $1/\phi\in\mathcal{H}^{\infty}(\mathbb{C}_{+})$, hence $\hat{\dot{y}}\in\mathcal{H}^{2}(\mathbb{C}_{+})$ (see §2.1 for the definition of Hardy spaces). From (12) we see that the description of $\mathcal{Y}(0,\infty)$ is equivalent, through the Laplace transform, to the description of these $F\in\mathcal{H}^{2}(\mathbb{C}_{+})$ such that $\phi F$ lies again in $\mathcal{H}^{2}(\mathbb{C}_{+})$. Proposition 6 below allows us to characterize such functions $F$, under a growth assumption on $\phi$.
@@
Let $\phi$ be an entire function of order111see Definition 19 <1<1. Then, for every $F\in\mathcal{H}^{2}(\mathbb{C}_{+})$, the product $\phi F$ lies in $\mathcal{H}^{2}(\mathbb{C}_{+})$ if and only if
@@
We will obtain Proposition 6 as a consequence of the following assertion:
$\operatorname{supp}\mathcal{F}\phi(i\cdot)\subset\{0\}.$ (13)
On the one hand, a function $\phi$ entire on $\mathbb{C}$ with order <1<1 may not even lie in $\mathcal{S}^{\prime}(\mathbb{R})$ (e.g. $\phi(s)=\cosh\sqrt{s}$), hence (13) is a formal statement. On the other hand, the Paley–Wiener theorem [rudin, Theorem 19.3, p. 370] asserts that if $f\in\mathcal{H}(\mathbb{C})\cap L^{2}(\mathbb{R})$ satisfies
@@
for some $A>0$, then $\operatorname{supp}\mathcal{F}f\subset[-A,A]$. As a function $\phi$ of order <1<1 satisfies the above estimate for any $A>0$, the above Paley-Wiener theorem is somehow compatible with (13). Lemma 20 will give a meaning to (13).
@@
$\|\varphi\|_{\mathcal{G}_{s,R,\gamma}}^{2}:=\sum_{n=0}^{\infty}\left(\frac{\|\varphi^{(n)}\|_{L^{2}(\mathbb{R})}}{M_{n}}\right)^{2}<\infty,\quad M_{n}:=\frac{(ns)!}{R^{ns}}(1+n)^{-s\gamma-1/4},$ (15)
where $(ns)!:=\Gamma(ns+1)$. The above defines a Hilbert space $\mathcal{G}_{s,R,\gamma}$, which is continuously and densely embedded in $L^{2}(\mathbb{R},dx)$.
@@
To our knowledge, Theorem 7 is new even for $\gamma=0$. Indeed, what is well-established is that a function $\varphi\in C_{c}^{\infty}(\mathbb{R})$ such that
$\exists R,C>0,\quad\forall k\in\mathbb{N},\quad\forall x\in\mathbb{R},\quad|\varphi^{(k)}(x)|\leq C\frac{(ns)!}{R^{ns}},$
@@
$\exists L,C^{\prime}>0,\quad\forall\xi\in\mathbb{R},\quad|\mathcal{F}\varphi(\xi)|\leq C^{\prime}e^{-L|\xi|^{1/s}},$
and the converse also holds (see e.g. [gs2, Chapter IV],[Rodino, Theorem 1.6.1]). In the previous works we are aware of, a link between $R$ and $L$ can be made explicit, but it is never sharp.
@@ 1.3 Consequences for the interpolation problem in Gevrey-$2$ class
We now consider the interpolation problem in Gevrey classes [^2] We refer to [bilodeau] for a historical survey. For more results on the interpolation problem, notably on more general Denjoy-Carleman classes, see [wahde, petzsche, kiro]. of order $s>1$. For fixed $s>1$, we consider a sequence $(a_{n})$ of complex numbers such that
$\exists R_{0},C>0,\quad\forall n\in\mathbb{N},\quad|a_{n}|\leq C\frac{(ns)!}{R_{0}^{ns}}.$ (16)
We say that a function $\varphi\in C^{\infty}(\mathbb{R})$ interpolates $(a_{n})$ (at $t=0$) if $\varphi^{(n)}(0)=a_{n}$ for every $n$. From Borel’s Theorem [^3] Borel published this result in 1895 within his thesis [Bobo], although it seems to have been proved first by Peano in 1884, see [borel]. we know that such a function $\varphi$ exists, regardless of the growth hypothesis made on $(a_{n})$. In [carleson] Carleson studies the universal moment problem and establishes a sufficient condition for it to be solved. The moment problem is connected with the interpolation problem through the formula
$\varphi^{(n)}(0)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\xi^{n}\mathcal{F}\varphi(-\xi)d\xi,\quad\mathcal{F}\varphi(\xi):=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-it\xi}\varphi(t)dt.$
As a consequence of his result on the moment problem, Carleson obtains that one may always interpolate a sequence $(a_{n})$ satisfying (16) by a function $\varphi$ which is Gevrey of order $s$, i.e. satisfying
$\exists R_{1},C^{\prime}>0,\quad\forall n\in\mathbb{N},\quad\forall t\in\mathbb{R},\quad|\varphi^{(n)}(t)|\leq C^{\prime}\frac{(ns)!}{R_{1}^{ns}}.$ (17)
For the applications in control theory we have in mind [reachable_martin, rosier, laurent], understanding how the parameter $R_{1}$ in (17) depends on the parameter $R_{0}$ in (16) is crucial in order to obtain sharp reachability results. Around the same year Carleson’s results are published, Mitjagin [mitjagin] independently shows that one can always take $R_{1}=\rho_{s}R_{0}-\epsilon$, where $\rho_{s}:=\cos(\pi/2s)<1$ and $\epsilon$ is any prescribed positive number. In [mitjagin], the interpolating function $\varphi$ is defined on the smaller interval [1,1][-1,1], which is not essential, as can be seen by cut-off arguments, see [reachable_martin, Lemma 3.7]. We shall refer to the numerical constant $\rho_{s}$ as the loss factor of the interpolation problem, as it prescribes how one should relax the growth assumptions when passing from (16) to (17). The constant $\rho_{s}$ was further shown by Mitjagin to be sharp in the sense that for every $\epsilon,R_{0}>0$ and $s>1$, there exists a sequence $(a_{n})$ satisfying (16) which can be interpolated by no function $\varphi$ satisfying (17) with $R_{1}\geq\rho_{s}R_{0}+\epsilon$. Mitjagin’s proof is quite elegant, and the optimality of the constant $\rho_{s}$ is deduced from its optimality in a Phragmén-Lindelöf principle. The paper [mitjagin] was originally written in Russian, and its translation is rather hard to access. Furthermore, in view of its size (4 pages), the published version seems to offer only a sketch of the proof rather than a fully detailed demonstration. Hence, for the sake of completeness, we give in Appendix A.2 a detailed proof of Mitjagin’s solution of the interpolation problem.
As a consequence of our study of the trackable subspace $\mathcal{Y}(0,T)$ we can make more precise the relation between $(a_{n})$ and $\varphi$, when $s=2$. The following result characterizes the range of the Borel operator $\mathcal{B}_{t_{0}}$, defined on smooth functions by $\mathcal{B}_{t_{0}}\varphi:=(\varphi^{(n)}(t_{0}))_{n=0}^{\infty}$, over Hilbert spaces of Gevrey functions. We will use the following notation: for an interval $I\subset\mathbb{R}$ and parameters $s,R>0$ and $\gamma\in\mathbb{R}$, we let $\mathcal{G}_{s,R,p}(I)$ be the Hilbert space defined similarly as in (15), with $L^{2}(I)$ in place of $L^{2}(\mathbb{R})$.
Proposition 9.
Let $I\subset\mathbb{R}$ be a non trivial interval and $t_{0}$ lying in its interior. Let also $R>0$ and $p\in\mathbb{Z}$. Then
and
In the above, $A^{2}(\Omega)$ stands for the Bergman space on $\Omega$, that is the set of functions which are holomorphic and square integrable on $\Omega$.
For an arbitrary sequence $(a_{n})$, we deduce a formula for the largest $R>0$ such that $(a_{n})$ can be interpolated by a function $\varphi$ belonging to the set $G^{2,R}[-1,1]$, which is defined similarly as in (3) with [1,1][-1,1] in place of $[0,T]$. To state the result we introduce the notation
$R_{a}:=\sup\{R>0:\exists\varphi\in G^{2,R}[-1,1],\quad\forall n\in\mathbb{N},\quad\varphi^{(n)}(0)=a_{n}\},$
where it is understood that $R_{a}=0$ if the above supremum runs over the empty set.
@@
The rest of the paper is organized as follows. In Section 2 we show our tracking results Theorems 1 and 2, as well as the intermediate results Proposition 6 and Theorem 7, and finally Proposition 4. In Section 3 we explain how our methods can be applied to systems close to (1), having different boundary conditions or smoother control laws, or where the output is given by the Dirichlet trace at a given point of $(0,1)$. In Section 4 we show Proposition 9 and Corollary 10. In Appendix A we study the interpolation problem in Gevrey classes of order $s>1$, notably by providing a detailed proof Mitjagin’s [mitjagin, Theorem 1 and 1a].